# Water-rich sub-Neptunes and rocky super Earths around different Stars: Radii shaped by Volatile Partitioning, Formation, and Evolution

### v1

@@ Abstract
Context. Despite precise characterization measurements, the nature of planets with radii ranging from 2 to 4 Earth radii, the sub-Neptunes, remains unknown due to degeneracies in interior models. However, a statistical ensemble of small planets with measured masses and radii around different stars has been compiled by the field.
@@
Results. Out of the tested scenarios, the mass-radius relation of exoplanets is best matched under the mixed assumption without water sequestration to the interior. We quantify the radius valley location and scaling with stellar mass. Fractionation is not found to significantly alter the composition of the planets for our initial conditions due to initially massive envelopes on all planets. In contrast, water sequestration has a profound effect on the radius evolution and compositional budget of the planets. The model predicts the preservation of large quantities of water even if the gaseous envelope is lost. Planets with corresponding bulk densities are not observed in comparably large numbers.
Conclusions. By combining formation and evolution model, we probe a parameter space favored by core accretion theory. We conclude that observationally searching for the imprint of fractionation on present-day exoplanet atmospheres can constrain initial gas mass fractions of exoplanets but not the occurrence of photoevaporative mass loss in general. The dissolution of different volatiles into the planetary interior and solidification of the magma ocean are natural next steps for comprehensive treatment of atmosphere-interior interaction.
@@
So far, the mass, orbital period, and radius of around two hundred sub-Saturn-mass exoplanets have been characterized with high precision (Parc et al., 2024). This success was possible thanks to various observational programs, such as K2 (Howell et al., 2014) or TESS (Ricker et al., 2015), which find nearby planets amenable to characterization via photometric and radial-velocity follow-up observations (Astudillo-Defru et al., 2020; Bonomo et al., 2023; Bonfanti et al., 2024; Brinkman et al., 2023; Damasso et al., 2023; Cloutier et al., 2024; Giacalone et al., 2022; Kossakowski et al., 2021; Leleu et al., 2024b; Luque & Pallé, 2022; Luque et al., 2022, 2023; Orell-Miquel et al., 2023; Osborn et al., 2021; Passegger et al., 2024; Ulmer-Moll et al., 2023). The set of exoplanet data complements the statistical study of exoplanet occurrence from Kepler (Borucki et al., 2010; Lissauer et al., 2023) where the masses of the planets are often unconstrained.
@@
But even with a precisely determined mass, the interior structure of the planet is not fully characterized. This is the case for rocky planets due to an a priori unknown metallic core to rocky mantle ratio (Dorn et al., 2015) and even more so if volatile species such as water, hydrogen (H), and helium (He) are present. Without additional constraints from atmospheric spectroscopy, this degeneracy is impossible to break for an individual object. However, the statistical study of the exoplanet demographics can reveal further insights into the typical structure of an average exoplanet. One key feature which was discovered is the so-called radius valley which separates the smaller super Earths from the larger sub-Neptunes (Fulton et al., 2017; Fulton & Petigura, 2018; Van Eylen et al., 2018). For planets with gaseous envelopes which are continuously stripped by photoevaporation, driven by high-energy radiation from the star, such a transition was predicted to exist (Owen & Wu, 2013; Lopez & Fortney, 2013; Jin et al., 2014; Chen & Rogers, 2016) and offered a compelling explanation. With the increasing knowledge of the mass and thus the bulk density of these planets, Luque & Pallé (2022) found that – at least for M-dwarf planets – the radius valley as seen by Kepler is fundamentally the projection of a gap in planetary bulk density, as expected from a compositional transition.
@@
An open question remains: what is the volatile material present on sub-Neptunes but absent on super Earths? Traditionally, it was thought to be pure H/He and large quantities of water ice were excluded (Lopez, 2017; Jin & Mordasini, 2018; Rogers & Owen, 2021). However, water accreted as ice likely changes its phase after accretion and could be present in large quantities as steam (partially) mixed with H/He (Dorn et al., 2018; Jin & Mordasini, 2018; Zeng et al., 2019; Mousis et al., 2020; Venturini et al., 2020; Turbet et al., 2020; Acuña et al., 2021; Aguichine et al., 2021; Pierrehumbert, 2023). Here, we will explore scenarios where water mixes with H/He, as well as a case with two distinct, stratified layers.
@@
If water accounts for a significant portion (on the order of tens of percent) of the total mass of the planet, it likely accreted it in solid form beyond the water iceline in the protoplanetary disk. Therefore, the presence or absence of water on sub-Neptunes is a fundamental probe for large scale migration of these planets (Lin et al., 1996; Ward, 1997; Paardekooper et al., 2023) and provides essential constraints for modern planet formation modeling. The orbits of sub-Neptune pairs are commonly in mean-motion resonance (Leleu et al., 2024a; Chen et al., 2024), which is likely linked to migration. However, it can be reproduced by small-scale convergent migration. Verifying the migration hypothesis through compositional analysis would prove that migration of sub-Neptunes over significant distances occurred.
@@
Indeed, the limited masses reached by purely rocky planets forming via core accretion strongly favors that the origin of sub-Neptune mass planets lies at larger distances (Alibert et al., 2005; Emsenhuber et al., 2023b). Therefore, a water-rich composition is the focus of studies coupling formation and evolution models to explore the radius valley (Venturini et al., 2020; Chakrabarty & Mulders, 2024; Burn et al., 2024).
@@
Additionally, early evidence from JWST transmission spectroscopy suggests that some of the observed sub-Neptune atmospheres are likely water-rich. So far, the presence of water was confirmed in TOI-270 d (Holmberg & Madhusudhan, 2024; Benneke et al., 2024), GJ 3470 b (Beatty et al., 2024), and GJ 9827 d (Piaulet-Ghorayeb et al., 2024), the latter begin consistent with a water-rich ”steam world”. Tentative detections have been obtained for other sub-Neptunes as well (MacDougall et al., 2023; Kempton et al., 2023; Wallack et al., 2024). Further observations are scheduled for this population, which have the potential to quantify the atmospheric water abundance (Kempton & Knutson, 2024; Nixon et al., 2024). Combined with data on the bulk properties, these measurements help to reveal the internal composition of sub-Neptunes.
@@
If the atmosphere consists of more than one species, it is possible that atmospheric escape preferentially leads to loss of the lighter species relative to the heavier species. If hydrogen is lost in larger abundance than the heavier species, this would change the metallicity in the envelope over time and could lead to a metal-rich late atmosphere. This process, called fractionation, was first studied analytically (Hunten et al., 1987; Zahnle & Kasting, 1986; Zahnle et al., 1990) and is a common subject in the context of Solar System planets’ atmospheres (Chassefière, 1996; Odert et al., 2018; Lammer et al., 2020). For exoplanets, the process was investigated for the fractionation of He from H (Hu et al., 2015; Malsky & Rogers, 2020; Cherubim et al., 2024) and found large He fractions in some cases. Gu & Chen (2023) reported significant deuterium fractionation on sub-Neptunes occurring in a part of the parameter space. So far, no studies address the fractionation of heavier elements in sub-Neptune atmospheres. In this work, we investigate using a simple analytic approach the fractionation of oxygen (from assumed dissociated, abundant water molecules) from H, which is a required addition for water-rich sub-Neptunes.
@@
In addition to the exact atmospheric partitioning, there is another large source of uncertainty on the water partitioning in the planet (see Ikoma et al., 2018; Lichtenberg et al., 2023, for recent reviews). Instead of mixing with the atmosphere, water can be sequestered in the deep interior (mantle and core). While solidified rocky mantle material can only hold limited amounts of volatiles (Shah et al., 2021), large quantities of volatiles can be dissolved in magma (Hirschmann et al., 2012; Dorn & Lichtenberg, 2021; Sossi et al., 2023). Moreover, during core formation, volatiles can also get sequestered in the iron core (Luo et al., 2024). At temperatures above 2000 K, supercritical rock and water can mix (Vazan et al., 2022). The presence of volatiles in the deeper interior decreases their amount in the envelope. In consequence, the total radius changes which will have interesting effects on the evolution of the planet, especially on photoevaporative mass loss which is sensitive to the planetary radius. Evolution models coupled to the interior were developed for the case of (mostly) hydrogen dissolution into the magma ocean by Chachan & Stevenson (2018); Kite et al. (2019, 2020); Lichtenberg (2021); Schlichting & Young (2022) who found significant impacts on the resulting radii. In particular, Kite et al. (2019) found a truncation of the radius distribution at the upper end of the sub-Neptune radii (the radius cliff) when considering the non-ideality of the hydrogen in contact with the magma which leads to more mass being dissolved. Recently, Rogers et al. (2024) developed an evolution model for super Earths including the effect of water dissolution into magma (Kite & Schaefer, 2021, see also) and iron core following Luo et al. (2024). They use it to infer allowed water mass fractions for super Earths and generally treat the initial conditions as free parameters. Here, complementary, we present a similar model but use initial conditions from formation modeling instead to predict the effect of water sequestration on the population of simulated planets.
@@
Another aspect of exoplanet demographics is the dependency on stellar properties, in particular the stellar mass. Several studies have focused on stellar mass dependency of the radius valley (Berger et al., 2020; Cloutier & Menou, 2020; Van Eylen et al., 2021; Petigura et al., 2022; Luque & Pallé, 2022; Ho et al., 2024; Bonfanti et al., 2024; Parc et al., 2024) and agree in inferring the locus of the valley at lower radii around smaller stars as predicted by theoretical works (Gupta & Schlichting, 2020; Rogers et al., 2023). This trend was tested including water-rich atmospheres in Venturini et al. (2024) using the prediction of a formation model where pebble accretion is the mode of solid mass delivery. They confirm a persistent density valley and report a pollution of the radius valley due to lower-mass water worlds around low-mass stars. Previously, scatter in these relations was often linked to collisions (Inamdar & Schlichting, 2015; Izidoro et al., 2022). Venturini et al. (2024) include the effect of collisions but rely on estimates using single-planet formation modeling. Here, we use results from full N-body interactions between 50 growing planets per disk. However, we use a simple treatment of collision outcomes, slightly modified compared to our preceding study focusing on Solar-mass stars (Burn et al., 2024).
@@ 2.1 Formation
The (exo-)planetary mass and radius distribution are shaped by both the embedded stage in the protoplanetary disk as well as the long-term evolution of each planet. For the former, which we call here the formation stage, we use the results of the population synthesis work by Burn et al. (2021). In that work, for five different stellar masses (0.1, 0.3, 0.5, 0.7, and 1.0 M⊙), an initial protoplanetary disk with 50 embedded, moon-mass embryos was assumed. Since the protoplanetary disk properties are expected to depend on the stellar mass, the initial disk masses following the observations of Class I disks by Tychoniec et al. (2018) were scaled linearly with the stellar mass. Similarly, the disk radius decreases for lighter disks (Andrews et al., 2010, 2018, ) while the disk lifetime, solid mass fraction, and inner edge orbital period were chosen to not depend on the stellar mass. In Emsenhuber et al. (2023a), it was confirmed to a large degree that such initial conditions can lead to the observed, evolved Class II disk properties. The approach to randomize those initial conditions allows for statistical comparison of the resulting planets with observations, which is the idea of population syntheses (Mordasini & Burn, 2024; Burn & Mordasini, 2024). The disk subsequently evolves, which is modeled in Burn et al. (2021) assuming a turbulent viscosity parameter of and a simple photoevaporation prescription tuned such that the disk lifetimes are distributed around 3 Myr for all modeled stars 111While there is a trend of shorter disk lifetimes around stars more massive than 1.5 M⊙ compared to those lighter than this threshold mass, this trend is not confirmed to exist below 1 M⊙ (Richert et al., 2018, see also the discussion in Burn et al., 2021 but note the exception of some long-lived disks Silverberg et al., 2020).
@@
The formation of planets as modeled by Burn et al. (2021) uses the model described in detail in Emsenhuber et al. (2021) which is based on the works of Alibert et al. (2005); Mordasini et al. (2009, 2012b). It includes the N-body forces between the embryos (using mercury, Chambers, 1999) and an analytically described accretion of planetesimals by a protoplanet with a gaseous envelope (Inaba et al., 2001; Inaba & Ikoma, 2003; Fortier et al., 2013), which dominate the first modeled stage of growth. At intermediate masses, the accretion of gas becomes important which is calculated by solving the one-dimensional internal structure equations (Bodenheimer & Pollack, 1986) of the gaseous envelope and tracking its energy content which allows the description of its cooling and contraction. As long as the planet’s growth is not limited by the mass of the surrounding gaseous disk, the amount of gas accreted is given by the contraction of the envelope. When the planet mass becomes dominated by gas instead of solids, the larger compressibility of gas allows for increased contraction which leads to even more gas accretion. Thus, the planet enters a runaway accretion stage (Makino et al., 1998). This rapid gas accretion phase is stopped once the disk is not able to supply by viscous transport the amount of gas that could be accreted by the planet given its cooling and contraction rate. At this point, the planet detaches from the gaseous disk, contracts, and its accretion rate is modeled using an analytical estimate for the maximum gas accretion rate (Emsenhuber et al., 2021).
@@
While the disk is present, it will be perturbed by the young planet and exert a torque on it. Therefore, the planets orbital elements are changed and in our model follows the prescriptions of Paardekooper et al. (2011) for type I migration and Dittkrist et al. (2014) for type II migration. Migration will have the effect of typically moving volatile-rich planets formed beyond the snow line toward the star. Therefore, a common outcome to models including planetary migration is the prediction of planets with volatile (mostly water) contents of several tens of percent in bulk mass at short orbital periods (Ida & Lin, 2008; Alibert et al., 2005). This effect is even more pronounced for lower stellar masses due to the lack of planets reaching the slower type II regime (Alibert & Benz, 2017; Burn et al., 2021). Nevertheless, in the simulations of Burn et al. (2021), which form the basis of our calculations, rocky planets emerging interior to the water iceline can also form due to an assumed high efficiency of planetesimal formation in this region. This is in contrast to other simulations assuming reduced efficiency of accretion of dry particles or different initial planetesimal distributions (Coleman et al., 2019; Miguel et al., 2020).

### v2

@@ Abstract
Context. Despite precise characterization measurements, the nature of planets with radii ranging from 2 to 4 Earth radii, the sub-Neptunes, remains unknown due to degeneracies in interior models. However, an impressive ensemble of small planets with measured masses and radii around different stars has been compiled by the field.
@@
Results. Out of the tested scenarios, the mass-radius relation of exoplanets is relatively well matched by all scenarios where the atmosphere is of mixed composition. The agreement depends on mass, with better consistency for the model without dissolution below 3 Earth masses and hints of the opposite at higher masses. Fractionation is not found to significantly alter the properties of the planets for our initial conditions due to initially massive envelopes on all planets. For all scenarios, we quantify the radius valley location and scaling with stellar mass.
Conclusions. Water-rich sub-Neptunes mass-radius relations are broadly consistent with observations but statistical surveys in mass and radius are required for distinction of the scenarios. The dissolution of different volatiles into the planetary interior and solidification of the magma ocean are natural next steps toward a comprehensive treatment of atmosphere-interior interaction in planet evolution models.
@@
So far, the mass, orbital period, and radius of around two hundred sub-Saturn-mass exoplanets have been characterized with high precision (Parc2024). This success was possible thanks to various observational programs, such as K2 (Howell2014) or TESS (Ricker2015), which find nearby planets amenable to characterization via photometric and radial-velocity follow-up observations (Astudillo-Defru2020; Bonomo2023; Bonfanti2024; Brinkman2023; Damasso2023a; Cloutier2024; Giacalone2022; Kossakowski2021; Leleu2024; Luque2022; Luque2022b; Luque2023; Orell-Miquel2023; Osborn2021; Passegger2024; Ulmer-Moll2023). The set of exoplanet data complements the statistical study of exoplanet occurrence from Kepler (Borucki2010; Lissauer2023) where the masses of the planets are often unconstrained.
@@
But even with a precisely determined mass, the interior structure of the planet is not fully characterized. This is the case for rocky planets due to an a priori unknown metallic core to rocky mantle ratio (Dorn2015) and even more so if volatile species such as water, hydrogen (H), and helium (He) are present. Without additional constraints from atmospheric spectroscopy, this degeneracy is impossible to break for an individual object. However, the statistical study of the exoplanet demographics can reveal further insights into the typical structure of an average exoplanet. One key feature which was discovered is the so-called radius valley which separates the smaller super Earths from the larger sub-Neptunes (Fulton2017; Fulton2018; VanEylen2018). For planets with gaseous envelopes which are continuously stripped by photoevaporation, driven by high-energy radiation from the star, such a transition was predicted to exist (Owen2013; Lopez2013; Jin2014; Chen2016) and offered a compelling explanation. With the increasing knowledge of the mass and thus the bulk density of these planets, Luque2022 found that – at least for M-dwarf planets – the radius valley as seen by Kepler is fundamentally the projection of a gap in planetary bulk density, as expected from a compositional transition.
@@
An open question remains: what is the volatile material present on sub-Neptunes but absent on super Earths? Traditionally, it was thought to be pure H/He and large quantities of water ice were excluded (Lopez2017; Jin2018; Rogers2021a). However, water accreted as ice likely changes its phase after accretion and could be present in large quantities as steam (partially) mixed with H/He (Dorn2018; Jin2018; Zeng2019; Mousis2020; Venturini2020; Turbet2020; Acuna2021; Aguichine2021; Pierrehumbert2023). Here, we will explore scenarios where water mixes with H/He, as well as a case with two distinct, stratified layers.
@@
If water accounts for a significant portion (on the order of tens of percent) of the total mass of the planet, it likely accreted it in solid form beyond the water iceline in the protoplanetary disk. Therefore, the presence or absence of water on sub-Neptunes is a fundamental probe for large scale migration of these planets (Lin1996; Ward1997; Paardekooper2023) and provides essential constraints for modern planet formation modeling. The orbits of sub-Neptune pairs are commonly in mean-motion resonance (Leleu2024a; Chen2024), which is likely linked to migration. However, it can be reproduced by small-scale convergent migration. Verifying the migration hypothesis through compositional analysis would prove that migration of sub-Neptunes over significant distances occurred.
@@
Indeed, the limited masses reached by purely rocky planets forming via core accretion strongly favors that the origin of sub-Neptune mass planets lies at larger distances (Alibert2005; Emsenhuber2023a). Therefore, a water-rich composition is the focus of studies coupling formation and evolution models to explore the radius valley (Venturini2020; Chakrabarty2024; Burn2024).
@@
Additionally, early evidence from JWST transmission spectroscopy suggests that some of the observed sub-Neptune atmospheres are likely water-rich. So far, the presence of water was confirmed in TOI-270 d (Holmberg2024; Benneke2024), GJ 3470 b (Beatty2024), and GJ 9827 d (Piaulet-Ghorayeb2024), the latter begin consistent with a water-rich ”steam world”. Tentative detections have been obtained for other sub-Neptunes as well (MacDougall2023; Kempton2023; Wallack2024). Further observations are scheduled for this population, which have the potential to quantify the atmospheric water abundance (Kempton2024; Nixon2024). Combined with data on the bulk properties, these measurements help to reveal the internal composition of sub-Neptunes.
@@
If the atmosphere consists of more than one species, it is possible that atmospheric escape preferentially leads to loss of the lighter species relative to the heavier species. If hydrogen is lost in larger abundance than the heavier species, this would change the metallicity in the envelope over time and could lead to a metal-rich late atmosphere. This process, called fractionation, was first studied analytically (Hunten1987; Zahnle1986; Zahnle1990) and is a common subject in the context of Solar System planets’ atmospheres (Chassefiere1996; Odert2018; Lammer2020). For exoplanets, the process was investigated for the fractionation of He from H (Hu2015; Malsky2020; Cherubim2024) and found large He fractions in some cases. Gu2023 reported significant deuterium fractionation on sub-Neptunes occurring in a part of the parameter space. So far, no studies address the fractionation of heavier elements in sub-Neptune atmospheres. In this work, we investigate using a simple analytic approach the fractionation of oxygen (from assumed dissociated, abundant water molecules) from H, which is a required addition for water-rich sub-Neptunes.
@@
In addition to the exact atmospheric partitioning, there is another large source of uncertainty on the water partitioning in the planet (see Ikoma2018; Lichtenberg2023, for recent reviews). Instead of mixing with the atmosphere, water can be sequestered in the deep interior (mantle and core). While solidified rocky mantle material can only hold limited amounts of volatiles (Shah2021), large quantities of volatiles can be dissolved in magma (Hirschmann2012; Dorn2021; Sossi2023). Moreover, during core formation, volatiles can also get sequestered in the iron core (Luo2024). At temperatures above 2000 K, supercritical rock and water can mix (Vazan2022). The presence of volatiles in the deeper interior decreases their amount in the envelope. In consequence, the total radius changes which will have interesting effects on the evolution of the planet, especially on photoevaporative mass loss which is sensitive to the planetary radius. Evolution models coupled to the interior were developed for the case of (mostly) hydrogen dissolution into the magma ocean by Chachan2018; Kite2019; Kite2020; Lichtenberg2021b; Schlichting2022 who found significant impacts on the resulting radii. In particular, Kite2019 found a truncation of the radius distribution at the upper end of the sub-Neptune radii (the radius cliff) when considering the non-ideality of the hydrogen in contact with the magma which leads to more mass being dissolved. Recently, Rogers2024a developed an evolution model for super Earths including the effect of water dissolution into magma (Kite2021, see also) and iron core following Luo2024. They use it to infer allowed water mass fractions for super Earths and generally treat the initial conditions as free parameters. Here, complementary, we present a similar model but use initial conditions from formation modeling instead to predict the effect of water sequestration on the population of simulated planets.
@@
Another aspect of exoplanet demographics is the dependency on stellar properties, in particular the stellar mass (but see also Chen2025, for the dependency on metallicity). Several studies have focused on stellar mass dependency of the radius valley (Berger2020a; Cloutier2020; VanEylen2021; Petigura2022; Luque2022; Ho2024; Bonfanti2024; Parc2024) and agree in inferring the locus of the valley at lower radii around smaller stars as predicted by theoretical works (Gupta2020; Rogers2023). This trend was tested including water-rich atmospheres in Venturini2024 using the prediction of a formation model where pebble accretion is the mode of solid mass delivery. They confirm a persistent density valley and report a pollution of the radius valley due to lower-mass water worlds around low-mass stars. Previously, scatter in these relations was often linked to collisions (Inamdar2015; Izidoro2022). Venturini2024 include the effect of collisions but rely on estimates using single-planet formation modeling. Here, we use results from full N-body interactions between 50 growing planets per disk. However, we use a simple treatment of collision outcomes, slightly modified compared to our preceding study focusing on Solar-mass stars (Burn2024).
@@ 2.1 Formation
The (exo-)planetary mass and radius distribution are shaped by both the embedded stage in the protoplanetary disk as well as the long-term evolution of each planet. For the former, which we call here the formation stage, we use the results of the population synthesis work by Burn2021. In that work, for five different stellar masses (0.1, 0.3, 0.5, 0.7, and 1.0 M⊙), an initial protoplanetary disk with 50 embedded, moon-mass embryos was assumed. Since the protoplanetary disk properties are expected to depend on the stellar mass, the initial disk masses following the observations of Class I disks by Tychoniec2018 were scaled linearly with the stellar mass. Similarly, the disk radius decreases for lighter disks (Andrews2010; Andrews2018, ) while the disk lifetime, solid mass fraction, and inner edge orbital period were chosen to not depend on the stellar mass. In Emsenhuber2023, it was confirmed to a large degree that such initial conditions can lead to the observed, evolved Class II disk properties. The approach to randomize those initial conditions allows for statistical comparison of the resulting planets with observations, which is the idea of population syntheses (Mordasini2024; Burn2024a). The disk subsequently evolves, which is modeled in Burn2021 assuming a turbulent viscosity parameter of and a simple photoevaporation prescription tuned such that the disk lifetimes are distributed around 3 Myr for all modeled stars 111While there is a trend of shorter disk lifetimes around stars more massive than 1.5 M⊙ compared to those lighter than this threshold mass, this trend is not confirmed to exist below 1 M⊙ (Richert2018, see also the discussion in Burn2021 but note the exception of some long-lived disks Silverberg2020).
@@
The formation of planets as modeled by Burn2021 uses the model described in detail in Emsenhuber2020a which is based on the works of Alibert2005; Mordasini2009; Mordasini2012c. It includes the N-body forces between the embryos (using mercury, Chambers1999) and an analytically described accretion of planetesimals by a protoplanet with a gaseous envelope (Inaba2001; Inaba2003; Fortier2013), which dominate the first modeled stage of growth. At intermediate masses, the accretion of gas becomes important which is calculated by solving the one-dimensional internal structure equations (Bodenheimer1986) of the gaseous envelope and tracking its energy content which allows the description of its cooling and contraction. As long as the planet’s growth is not limited by the mass of the surrounding gaseous disk, the amount of gas accreted is given by the contraction of the envelope. When the planet mass becomes dominated by gas instead of solids, the larger compressibility of gas allows for increased contraction which leads to even more gas accretion. Thus, the planet enters a runaway accretion stage (Makino1998). This rapid gas accretion phase is stopped once the disk is not able to supply by viscous transport the amount of gas that could be accreted by the planet given its cooling and contraction rate. At this point, the planet detaches from the gaseous disk, contracts, and its accretion rate is modeled using an analytical estimate for the maximum gas accretion rate (Emsenhuber2020a).
@@
While the disk is present, it will be perturbed by the young planet and exert a torque on it. Therefore, the planets orbital elements are changed and in our model follows the prescriptions of Paardekooper2011 for type I migration and Dittkrist2014 for type II migration. Migration will have the effect of typically moving volatile-rich planets formed beyond the snow line toward the star. Therefore, a common outcome to models including planetary migration is the prediction of planets with volatile (mostly water) contents of several tens of percent in bulk mass at short orbital periods (Ida2008; Alibert2005). This effect is even more pronounced for lower stellar masses due to the lack of planets reaching the slower type II regime and shorter distance from the snowline to the inner system (Alibert2017; Burn2021). Nevertheless, in the simulations of Burn2021, which form the basis of our calculations, rocky planets emerging interior to the water iceline can also form due to an assumed high efficiency of planetesimal formation in this region. This is in contrast to other simulations assuming reduced efficiency of accretion of dry particles or different initial planetesimal distributions (Coleman2019; Miguel2020).